OpenStudy (anonymous):

solve the system of equations by graphing x+y=2 y-x=2

hero (hero):

y = 2 - x y = 2 + x Graph those: 2 - x = 2+x 0 = 2x x = 0 y = 2

OpenStudy (anonymous):

wait what are the points?

OpenStudy (anonymous):

what are the two points to plot?

hero (hero):

I guess you were supposed to graph it on paper instead of on a graphing calc